# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ 3.2 Oracle Inequalities
For $\boldsymbol{\beta}\in\mathbb{R}^{n_{T}+p_{y}}$ and any subset $\tilde{S}\subseteq\mathopen{}\mathclose{{}\left\{1,\dots,n_{T}+p_{y}}\right\}$ with cardinality $\tilde{s}$, let $\boldsymbol{\beta}_{\tilde{S}}\in\mathbb{R}^{\tilde{S}}$ and $\boldsymbol{\beta}_{\tilde{S}^{c}}\in\mathbb{R}^{\tilde{S}^{c}}$. Define the compatibility constant
$\gamma_{w}^{2}=\underset{\tilde{S}\subseteq\{1,\dots,n_{T}+p_{y}\}}{\text{min}}\ \ \underset{{||\boldsymbol{\beta}_{\tilde{S}^{c}}||_{1}\leq 3||\boldsymbol{\beta}_{\tilde{S}}||_{1};\ \boldsymbol{\beta}\in\mathbb{R}^{n_{T}+p_{y}}\backslash\{0\}}}{\text{min}}\ \ \frac{\boldsymbol{\beta}^{\prime}\mathbf{W}\mathbf{W}^{\prime}\boldsymbol{\beta}}{T\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\boldsymbol{\beta}_{\tilde{S}}}\right|}\right|_{2}^{2}}\ \ ,$ (6)
and assume that $\gamma_{w}^{2}>0$. This implies that $\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\boldsymbol{\beta}_{\tilde{S}}}\right|}\right|_{1}^{2}\leq\tilde{s}\ \frac{\boldsymbol{\beta}^{\prime}\mathbf{W}\mathbf{W}^{\prime}\boldsymbol{\beta}}{T\gamma_{w}^{2}}.$
Assumption 5, named restricted eigenvalue condition (Bickel et al., 2009), restricts the smallest eigenvalue of $\frac{1}{T}\mathbf{W}\mathbf{W}^{\prime}$ as a function of sparsity. In particular, instead of minimizing over all of $\mathbb{R}^{n_{T}+p_{y}}$, the minimum in (6) is restricted to those vectors which satisfy $||\boldsymbol{\beta}_{\tilde{S}^{c}}||_{1}\leq 3||\widehat{\boldsymbol{\beta}}_{\tilde{S}}||_{1}$, where $\tilde{S}$ has cardinality $\tilde{s}$. Therefore, the compatibility constant is a lower bound for the restricted eigenvalue of the matrix $\frac{1}{T}\mathbf{WW}^{\prime}$ (Bühlmann and van de Geer 2011, p. 106). Note also that if $(n_{T}+p_{y})<T$, the restricted eigenvalue condition is trivially satisfied if $\frac{1}{T}\mathbf{WW}^{\prime}$ is nonsingular, since $\boldsymbol{\beta}_{\widehat{S}}^{\prime}\boldsymbol{\beta}_{\widehat{S}}\leq\boldsymbol{\beta}^{\prime}\boldsymbol{\beta}$ for every $\boldsymbol{\beta}\in\mathbb{R}^{n_{T}+p_{y}}$, and so $\frac{\boldsymbol{\beta}^{\prime}\mathbf{WW}^{\prime}\boldsymbol{\beta}}{T\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\boldsymbol{\beta}_{\widehat{S}}}\right|}\right|_{2}^{2}}\geq\frac{\boldsymbol{\beta}^{\prime}\mathbf{WW}^{\prime}\boldsymbol{\beta}}{T\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\boldsymbol{\beta}}\right|}\right|_{2}^{2}}\geq\underset{\boldsymbol{\beta}\in\mathbb{R}^{n_{T}+p_{y}}\setminus\{0\}}{\text{min}}\frac{\boldsymbol{\beta}^{\prime}\mathbf{WW}^{\prime}\boldsymbol{\beta}}{T\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\boldsymbol{\beta}}\right|}\right|_{2}^{2}}>0$. In other words, we require $\widehat{\psi}_{min}^{w}>0$, where $\widehat{\psi}_{min}^{w}$ is the minimum eigenvalue of $\frac{1}{T}\mathbf{WW}^{\prime}$.
@@
Let $\gamma_{x}^{2}$ be the compatibility constant of the restricted eigenvalue of $\frac{1}{T}\mathbf{XX}^{\prime}$. Since this captures how strongly predictors are correlated in the sample, as a consequence of the theoretical treatment of Section 2.1, we have $\gamma_{w}^{2}>\gamma_{x}^{2}$ with high probability as the degree of serial correlation increases (when both $\frac{1}{T}\mathbf{WW}^{\prime}$ and $\frac{1}{T}\mathbf{XX}^{\prime}$ are nonsingular, we have $\widehat{\psi}_{min}^{w}>\widehat{\psi}_{min}^{x}$ with high probability). Of course, $\gamma_{w}^{2}$ and $\gamma_{x}^{2}$ also depend on the cardinalities $\tilde{s}$ and $s$. However, here we emphasize the role of serial correlation.
@@
Let Assumptions 1, 34 and 5 hold. Furthermore, let the conditions of Theorem 4 hold. When $T$ and $n_{T}$ are sufficiently large. the following oracle inequalities hold simultaneously with probability at least $1-C\mathopen{}\mathclose{{}\left(\sqrt{\text{log(T)}}}\right)^{-1}$: ($a$) $\frac{1}{T}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\mathbf{W}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}\right|}\right|_{2}^{2}\leq\frac{4\tilde{s}\lambda^{2}}{\gamma_{w}^{2}}$; ($b$) $\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}}\right|}\right|_{1}\leq\frac{4\tilde{s}\lambda}{\gamma_{w}^{2}}$.
Proof: see Supplement B.5
Remark 3:
Under the additional assumption that $\tilde{s}\lambda\rightarrow 0$ one can also establish, as an immediate corollary to Theorem 5, the following convergence rates: ($a$) $\frac{1}{T}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\mathbf{W}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}\right|}\right|_{2}^{2}=O_{P}\mathopen{}\mathclose{{}\left(\frac{\tilde{s}}{T}\mathopen{}\mathclose{{}\left((n_{T}+p_{y})\mathopen{}\mathclose{{}\left(\sqrt{log(T)}}\right)}\right)^{2/c_{1}}}\right)$; ($b$) $\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}}\right|}\right|_{1}=O_{P}\mathopen{}\mathclose{{}\left(\frac{\tilde{s}}{\sqrt{T}}\mathopen{}\mathclose{{}\left((n_{T}+p_{y})\mathopen{}\mathclose{{}\left(\sqrt{log(T)}}\right)}\right)^{1/c_{1}}}\right)$.
3.3 Comparison with ARDL and GLS Estimators
The ARDL consists of regressing the response on its past realizations – the autoregressive component – as well as on current and past values of the predictors – the distributed lag component (see, e.g.,  Panopoulou and Pittis, 2004). Although this method does mitigate serial correlation, it has the drawback of requiring a very large number of coefficients to be estimated. This issue becomes particularly relevant when the sample size is small. In contrast, our proposal only requires the addition of a few response lags.
The popular Cochrane-Orcutt GLS estimator approximates the serial correlation structure of the error term while retaining consistent coefficient estimation (see, e.g.,  Cochrane and Orcutt, 1949). Although this improves statistical efficiency and inference compared to conventional least squares, it does not tackle directly the risk of spurious correlations due to predictors’ serial correlations, as described in Section 2.1. In particular, while the GLS filter may reduce predictors’ serial correlations, it does not remove them completely if the AR structure of the error term differs from the AR or ARMA structures of the predictors. Furthermore, compared to the ARMAr-LASSO, the GLS requires the additional step of estimating the error term $\varepsilon_{t}$. A detailed theoretical comparison between the GLS-LASSO of (Chronopoulos et al., 2023) and our ARMAr-LASSO is given in Supplement D. In Section 4 we will show that ARMAr-LASSO outperforms both ARDL-LASSO and GLS-LASSO in a variety of simulated scenarios and on real-world data.
@@ 4.1 Simulation Experiments
    General AR/ARMA Setting: ${x}_{j,t}=0.8{x}_{j,t-1}+u_{j,t}$; ${x}_{h,t}=0.6{x}_{h,t-1}+0.3{x}_{h,t-2}+u_{h,t}$; $x_{w,t}=0.5x_{w,t-1}+0.4x_{w,t-2}+u_{w,t}+0.3u_{w,t-1}$; $x_{k,t}=0.7x_{k,t-1}+u_{k,t}+0.4u_{k,t-1}$, for $t=1,\dots,T$, and $j=1,\dots,4;\ h=5,\dots,7;\ w=7,\dots,10;\ k=11,\dots,n$. The error terms are generated as $\varepsilon_{t}=0.7\varepsilon_{t-1}+0.2\varepsilon_{t-2}+\omega_{t}$.
  • (D)
    General AR/ARMA Setting with Common Factor: $x_{i,t}=f_{t}+z_{i,t}$, where $f_{t}=0.9f_{t-1}+\delta_{t}$. Idiosyncratic components and the error terms are generated as in (C).
@@
The shocks are generated as follows: $u_{i,t}\sim i.i.d.\ N(0,1)$ with $\mathopen{}\mathclose{{}\left(\mathbf{C}_{u}}\right){ij}=c_{ij}^{u}=\rho^{|i-j|}$, $\delta_{t},\ \eta_{i,t}\sim i.i.d.\ N(0,1)$ with $\mathopen{}\mathclose{{}\left(\mathbf{C}_{\eta}}\right){ij}=c_{ij}^{\eta}=\rho^{|i-j|}$, and $\omega_{t}\sim i.i.d.\ N(0,\sigma_{\omega}^{2})$. For the DGPs in (A) and (C) we set $\rho=0.8$, while for the DGPs in (B) and (D) we set $\rho=0.4$ to generate predictors primarily influenced by the common factor, with weakly correlated AR and/or ARMA idiosyncratic components. Finally, we vary the value of $\sigma_{\omega}^{2}$ to explore different signal-to-noise ratios (SNRs).
We compare our ARMAr-LASSO (ARMAr-LAS) with the standard LASSO applied to the observed time series (LAS), LASSO applied to the observed time series plus lags of $y_{t}$ (LASy), GLS-LASSO as proposed by Chronopoulos et al. (2023) (GLS-LAS), autoregressive distributed lag LASSO (ARDL-LAS), and FarmSelect as proposed by Fan et al. (2020), which employs LASSO on factor model residuals (FaSel). The performance of each method is evaluated based on average results from 1000 independent simulations, focusing on the coefficient estimation error (CoEr) obtained as $||\widehat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}||_{2}$, the Root Mean Square Forecast Error (RMSFE), and the percentages of true positives (%TP) and false positives (%FP) in selecting relevant predictors. Simulations have varying numbers of predictors (dimensionality), $n=50,150,300$, and a fixed sample size, $T=150$. In this way we cover low ($n=50$), intermediate ($n=150$), and high ($n=300$) dimensional scenarios. For all methods, the tuning parameter $\lambda$ is selected using the Bayesian Information Criterion (BIC). Finally, regardless of the choice of $n$, $\boldsymbol{\alpha}^{*}$ is always taken to have the first $10$ entries equal to $1$ and all others equal to 00. In this way, as $n$ varies, we also cover different levels of sparsity. In addition to the results presented below, we provide simulations with a much larger sample size $T$ in Supplement G, and simulations where our ARMAr-LASSO misspecifies the autoregressive model of the predictors in Supplement H.
4.1.1 Common AR(1) Restriction:
For DGPs (A) and (B) we investigate settings with different $\phi$ ($0.3,0.6,0.9,0.95$) and different SNR ($0.5,1,5,10$). For GLS-LAS, we estimate an AR(1) model on $\widehat{\varepsilon}_{t}$ (see Supplement D) and use the resulting autoregressive coefficient to filter both response and predictors. For ARDL-LAS, we consider one lag of the response and one lag of each predictor as additional regressors – bringing the number of terms undergoing selection to $n\times 2+1$. For the working model underlying ARMAr-LAS, the $\widehat{u}$’s are obtained by filtering each series with an AR(1) process, and we consider $p_{y}=1$; that is, we take one lag of $y_{t}$ as additional predictor. Results are presented in Table 1 for SNR values of 1 and 10 (complete results are provided in Supplement E). For each SNR, CoEr and RMSFE (both expressed in relative terms to the values obtained by LAS), as well as %TP, and %FP are given for every $n$ and $\phi$ considered (the best CoEr and RMSFE are in bold). Results, which are similar for the two DGPs, have ARMAr-LAS as the best performer in terms of CoEr and RMSFE across values of $\phi$, $n$ and SNR – demonstrating superior accuracy in both estimation and forecasting compared to the other LASSO-based methods considered. ARMAr-LAS also shows superior performance in feature selection, with a higher %TP and a lower %FP. These gains are more evident when serial correlations are stronger ($\phi=0.6$ or higher).
Notably, under the common AR(1) restriction, the ARMAr-LAS and GLS-LAS estimators should be equivalent (this is the one case where the GLS-LAS estimator removes the serial correlations of the predictors). Nevertheless, GLS-LAS performs on par with ARMAr-LAS only when serial correlations are low; ARMAr-LAS outperforms GLS-LAS when serial correlations are medium/high, likely because the latter requires the estimation of $\widehat{\varepsilon}_{t}$ (see Supplement D). Also, in some instances, ARDL-LAS exhibits a slightly lower %FP than ARMAr-LAS. However, this metric is calculated on $n\times 2+1$ predictors for the former; in terms of the absolute number of false positives, ARDL-LAS has more than ARMAr-LAS. Finally, we note that the superior performance of ARMAr-LAS in DGP (B) indicates its effectiveness in handling factor structures, where multicollinearities are more complex than for simple AR processes (DGP (A)).
4.1.2 General AR/ARMA Setting
@@
For DGPs (C) and (D) we consider $SNR=1,10$. These DPGs, where the common AR(1) restriction does not hold, represent more realistic scenarios. For GLS-LAS, we filter both response and predictors using the coefficients of an AR($p_{\varepsilon}$) model applied to $\widehat{\varepsilon}_{t}$, with the order $p_{\varepsilon}$ (max 2) selected with BIC. For ARDL-LAS, we consider two lags of the response and two lags of each predictor as additional regressors – bringing the number of term undergoing selection to $n\times 3+2$. For the working model underlying ARMAr-LAS, the $\widehat{u}$’s are obtained by filtering each series with an ARMA($p_{i},q_{i}$) process, with the orders $p_{i}$ and $q_{i}$ (max 2) selected via BIC. We consider $p_{y}=3$; that is, three lags of $y_{t}$ as additional predictors. Results are presented in Table 2. ARMAr-LAS outperforms all other LASSO-based methods in terms of estimation accuracy, forecasting, and feature selection in both DPG (C) and DPG (D), except for coefficients estimation (CoEr) when $SNR=1$, where ARDL-LAS performs slightly better. The effectiveness of our proposal in these more realistic settings highlights its suitability also when tackling differing AR and ARMA processes and common factors, where the working model (4) does not coincide with the true DGP of $y_{t}$.
@@
In Supplement F we compare the minimum eigenvalues of the predictors correlation matrix of ARMAr-LASSO with those of LASSO and GLS-LASSO. Results show that, for the four DGPs considered, ARMAr-LASSO relies on a correlation matrix which exhibits a larger minimum eigenvalue compared to those of the classical LASSO and GLS-LASSO.

### v3

@@ 3.4 ARMAr-LASSO: Oracle Inequalities
Consider $\mathbf{q}_{t}=(\widehat{\mathbf{w}}_{t}^{\prime},v_{t})^{\prime}$. There exist constants $c_{2}>c_{1}>2$ such that $\underset{i\leq n_{T}+p_{y}+1,\ t\leq T}{\text{max}}E(\mathopen{}\mathclose{{\left|q_{i,t}}}\right|^{2c_{2}})\leq C$ and $\underset{i\leq n_{T},t\leq T}{\text{max}}E(\mathopen{}\mathclose{{\left|x_{i,t}}}\right|^{2c_{2}})\leq C$.
@@
The error term $v_{t}$ is modeled as a stationary Near-Epoch-Dependent (NED) process which, under appropriate decay conditions, can be approximated by a strongly mixing sequence. In particular, this includes the case of stationary, and finite-order ARMA processes, which are well known to be strongly mixing with geometric decay of the mixing coefficients (see, e.g., davidson1994). For valid asymptotic results, we assume that the error term has finite $q$-th moments for some $q>4$, ensuring tails sufficiently light for our limit theory, although stronger assumptions such as exponential moment bounds could also be accommodated.
To derive the error bound of the ARMAr-LASSO estimator from (2), we follow the typical procedure presented in technical textbooks (see, e.g., Buhlmann2011, ch. 6). We need $\lambda$ to be sufficiently large as to exceed the empirical process $\underset{i\leq n_{T}+p_{y},t\leq T}{\text{max}}\mathopen{}\mathclose{{\left|\sum_{t=1}^{T}w_{i,t}v_{t}}}\right|$ with high probability.
Theorem 3:
Let Assumption 1 and 2 hold and define $\mathcal{A}_{T}\coloneqq\mathopen{}\mathclose{{\left\{\underset{i\leq n_{T}+p_{y},l\leq T}{\text{max}}\mathopen{}\mathclose{{\left|\sum_{t=1}^{l}\widehat{w}_{i,t}v_{t}}}\right|\leq\frac{T\lambda}{4}}}\right\}$. Furthermore, assume that $T$ and $n_{T}$ are sufficiently large as to have $\lambda\geq C\mathopen{}\mathclose{{\left(\sqrt{log(T)}}}\right)^{1/c_{1}}\frac{(2n_{T}+p_{y})^{1/c_{1}}}{\sqrt{T}}$. Then $\Pr(\mathcal{A}_{T})\geq 1-C\mathopen{}\mathclose{{\left(\sqrt{\text{log(T)}}}}\right)^{-1}$.
Theorem 3 establishes that the inequalities we need for the error bound of the proposed ARMAr-LASSO estimator hold with high probability. The bounds used in the proof of Theorem 3 put implicit limits on the divergence rate of $n_{T}$ relative to $T$. The term $\sqrt{\text{log($T$)}}$ is chosen arbitrarily as a sequence that grows slowly as $T\rightarrow\infty$. However, we can use any sequence that tends to infinity sufficiently slowly. For example, ADAMEK2023 use $\text{log}(\text{log}(T))$ to derive properties of the LASSO in a high-dimensional time series model under weak sparsity. We introduce the assumption on the “restricted” positive definiteness of the covariance matrix of the predictors, which allows us to generalize subsequent results to the high-dimensional framework.
Remark 5:
The concentration arguments in Theorem 3 are based on the NED-mixingale framework of ADAMEK2023, which allows for very general forms of temporal dependence, random regressors, and only finite moment assumptions. This choice is particularly suited to our projection-based and potentially misspecified regression setting with lagged dependent variables and estimated components. Sharper deviation bounds, and hence more aggressive tuning rates, could be obtained under stronger structural assumptions using, for example, the functional dependence approach of WuWu2016 or Fuk-Nagaev type inequalities for $\tau$- or $s$-mixing processes as in Babii2024. These alternatives, however, require additional restrictions on mixing rates, tail behavior, and the design structure. We therefore favor the more general framework of ADAMEK2023, which yields slightly more conservative but broadly valid theoretical guarantees.
Assumption 3:
[…]
Remark 6:
Let $\gamma_{x}^{2}$ be the compatibility constant of the restricted eigenvalue of $\frac{1}{T}\mathbf{XX}^{\prime}$. Since this captures how strongly predictors are correlated in the sample, as a consequence of the theoretical treatment of Sections 2.1 and 2.2, we have $\gamma_{\widehat{w}}^{2}>\gamma_{x}^{2}$ with high probability as the degree of serial correlation increases (when both $\frac{1}{T}\widehat{\mathbf{W}}\widehat{\mathbf{W}}^{\prime}$ and $\frac{1}{T}\mathbf{XX}^{\prime}$ are nonsingular, we have $\widehat{\psi}_{min}^{\widehat{w}}>\widehat{\psi}_{min}^{x}$ with high probability). Of course, $\gamma_{\widehat{w}}^{2}$ and $\gamma_{x}^{2}$ also depend on the cardinalities $\tilde{s}$ and $s$. However, here we emphasize the role of serial correlation.
@@
Let Assumptions 1, 2 and 3 hold. Furthermore, let the conditions of Theorem 3 hold. When assume that $T$ and $n_{T}$ are sufficiently large, the following oracle inequalities hold simultaneously with probability at least $1-C\mathopen{}\mathclose{{\left(\sqrt{\text{log(T)}}}}\right)^{-1}$: ($a$) $\frac{1}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\mathbf{W}}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}}\right|}}\right|_{2}^{2}\leq\frac{4\tilde{s}\lambda^{2}}{\gamma_{\widehat{w}}^{2}}$; ($b$) $\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}}}\right|}}\right|_{1}\leq\frac{4\tilde{s}\lambda}{\gamma_{\widehat{w}}^{2}}$. In addition, if $\underset{j\in\tilde{S}}{\text{min}}|\beta_{j}^{*}|>\frac{4\tilde{s}\lambda}{\gamma_{\widehat{w}}^{2}},$ ARMAr-LAS enjoys the variable screening property; that is, it correctly identifies all true non-zero coefficients.
Corollary 1:
Under the additional assumption that $\tilde{s}\lambda\rightarrow 0$ one can also establish, as an immediate corollary to Theorem 4, the following convergence rates: ($a$) $\frac{1}{T}\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\mathbf{W}}^{\prime}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*})}}\right|}}\right|_{2}^{2}=O_{P}\mathopen{}\mathclose{{\left(\frac{\tilde{s}}{T}\mathopen{}\mathclose{{\left((n_{T}+p_{y})\mathopen{}\mathclose{{\left(\sqrt{log(T)}}}\right)}}\right)^{2/c_{1}}}}\right)$; ($b$) $\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}}}\right|}}\right|_{1}=O_{P}\mathopen{}\mathclose{{\left(\frac{\tilde{s}}{\sqrt{T}}\mathopen{}\mathclose{{\left((n_{T}+p_{y})\mathopen{}\mathclose{{\left(\sqrt{log(T)}}}\right)}}\right)^{1/c_{1}}}}\right)$.
@@ 4.1 Simulation Experiments
    Common AR(2) restriction: $x_{i,t}=0.45x_{i,t-1}+0.45x_{i,t-2}+u_{i,t},\ \varepsilon_{t}=0.45\varepsilon_{t-1}+0.45\varepsilon_{t-2}+\omega_{t}$.
@@
    General AR/ARMA: $x_{i,t}=qf_{t}+z_{i,t}$, where $f_{t}=0.9f_{t-1}+\delta_{t}$, ${z}_{j,t}=0.8{z}_{j,t-1}+\eta_{j,t}$; ${z}_{h,t}=0.6{z}_{h,t-1}+0.3{z}_{h,t-2}+\eta_{h,t}$; $z_{w,t}=0.5z_{w,t-1}+0.4z_{w,t-2}+\eta_{w,t}+0.3\eta_{w,t-1}$; $z_{k,t}=0.7z_{k,t-1}+\eta_{k,t}+0.4\eta_{k,t-1}$, for $t=1,\dots,T$, and $j=1,\dots,4;\ h=5,\dots,7;\ w=7,\dots,10;\ k=11,\dots,n$. The error terms are generated as $\varepsilon_{t}=0.7\varepsilon_{t-1}+0.2\varepsilon_{t-2}+\omega_{t}$.
@@
The shocks are generated as follows: $u_{i,t}\sim i.i.d.\ N(0,1)$ with $\mathopen{}\mathclose{{\left(\mathbf{C}_{u}}}\right){ij}=c_{ij}^{u}=\rho^{|i-j|}$, $\delta_{t},\ \eta_{i,t}\sim i.i.d.\ N(0,1)$ with $\mathopen{}\mathclose{{\left(\mathbf{C}_{\eta}}}\right){ij}=c_{ij}^{\eta}=\rho^{|i-j|}$, and $\omega_{t}\sim i.i.d.\ N(0,\sigma_{\omega}^{2})$. For the DGP (A) and for the DGP (B) with $q=0$ we set $\rho=0.8$, while for the DGP (B) with $q=1$ we set $\rho=0.4$ to generate predictors primarily influenced by the common factor, with weakly correlated AR or ARMA idiosyncratic components. Finally, we vary the value of $\sigma_{\omega}^{2}$ to explore different signal-to-noise ratios (SNRs). For each DGPs, the performance of ARMAr-LASSO and benchmarks is evaluated based on average results from 1000 independent simulations, focusing on the coefficient estimation error (CoEr) obtained as $\widehat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}||_{2}$, the Root Mean Square Forecast Error (RMSFE), and the percentages of true positives (%TP) and false positives (%FP) in selecting relevant predictors. Regardless of the choice of $n$, $\boldsymbol{\alpha}^{*}$ is always taken to have the first $10$ entries equal to $1$ and all others equal to 0.
4.1.1 DGP(A): Filters Evaluation Under Common AR restriction.
We test several ARMAr-LASSO settings in terms of the number of lags of $y_{t}$ included in the model ($p_{y}$), the order of the AR filter used to obtain the estimated $\widehat{u}$’s ($p_{i}$), and the model selection method for the filter. We consider as reference setting the case where $p_{y}=3$ and the $\widehat{u}$’s are obtained by filtering each variable with an AR($p_{i}$), where $p_{i}$ (max 3) is selected via BIC and the AR parameters are estimated via ML (see Section 3.1). The dimensionality is kept fixed at $n=150$, while the sample size varies; we consider $T=75,150,300$. Both CoEr and RMSFE are relative to the working model 1, which includes the true $u$’s. Results are summarized in Table 1. To assess the impact of including lags of $y_{t}$, we exclude them from the penalization process and compare the cases with one lag ($y_{t-1}$) and three lags ($y_{t-1}^{t-3}$) treated as fixed predictors in the model. Note that under DGP (A), the number of lags of $y_{t}$ that yields white-noise error terms is $p_{y}=2$. Using too few lags ($y_{t-1}$) reduces both estimation and forecast accuracy, whereas using an excessive number of lags ($y_{t-1}^{t-3}$) does not provide additional benefits since, especially for large $T$, the performance coincides with that of ARMAr-LS. To evaluate the effect of filter misspecification, instead of selecting $p_{i}$ using BIC, we filter the variables with AR(1), AR(2), and AR(3) processes, thus under-specifying, correctly-specifying, and over-specifying the true order, respectively. Underestimating the true order worsens coefficient estimation and forecast accuracy, particularly at high SNRs, while over-specifying the true order provides no gains. Finally, we evaluate alternative model selection criteria other than BIC. In particular, we consider the Akaike Information Criterion (AIC), its small-sample corrected version (AICc), and block cross-validation (blCV), which is a time-series variant of cross-validation (racine1997). For blCV, we use a rolling window of size T3T-3, and for each variable, we select the order $p_{i}$ (max 3) that best predicts one step over the three out-of-sample horizons. All methods yield similar results, except for blCV, which performs worse than the others for high SNRs and large $T$.
4.1.2 DGP(B): General AR/ARMA.
@@
We compare our ARMAr-LASSO (ARMAr-LAS) with the standard LASSO applied to the observed time series (LAS), LASSO applied to the observed time series plus lags of $y_{t}$ (LASy), GLS-LASSO as proposed by chronopoulos2023 (GLS-LAS), autoregressive distributed lag LASSO (ARDL-LAS), and FarmSelect as proposed by Fan2020, which employs LASSO on factor model residuals (FaSel). For all methods, the tuning parameter $\lambda$ is selected using the Bayesian Information Criterion (BIC). For GLS-LAS, we filter both response and predictors using the coefficients of an AR($p_{\varepsilon}$) model applied to $\widehat{\varepsilon}_{t}$, with the order $p_{\varepsilon}$ (max 2) selected with BIC. For ARDL-LAS, we consider two lags of the response and two lags of each predictor as additional regressors – bringing the number of term undergoing selection to $n\times 3+2$. For the working model underlying ARMAr-LAS, the $\widehat{u}$’s are obtained by filtering each series with an ARMA($p_{i},q_{i}$) process, with the orders $p_{i}$ and $q_{i}$ (max 2) selected via BIC. We consider $p_{y}=3$; that is, three lags of $y_{t}$ as additional predictors. Simulations have varying numbers of predictors (dimensionality), $n=75,150,300$, and a fixed sample size, $T=150$. In this way, we cover low ($n=75$), intermediate ($n=150$), and high ($n=300$) dimensional scenarios, and also cover different levels of sparsity, consequently to $\boldsymbol{\alpha}^{*}||_{0}=10$. Results are presented in Table 2. ARMAr-LAS outperforms all other LASSO-based methods in terms of estimation accuracy, forecasting, and feature selection, regardless of the SNR and the presence of a common factor (i.e., when $q=1$). In particular, ARMAr-LASSO removes serial correlation in the predictors and the resulting spurious correlations, yielding more accurate estimation and forecasts than GLS-LASSO. Unlike GLS-LASSO, which fully eliminates predictor serial correlation only under the restrictive common AR($p$) condition, ARMAr-LASSO achieves this goal without requiring the same dynamic structure for predictors and errors. Moreover, ARMAr-LASSO requires only a few lags of $y_{t}$ as additional predictors, making it considerably more parsimonious than ARDL-LASSO, which quickly becomes over-parameterized when multiple lags are included. The effectiveness of our proposal in this realistic setting highlights its suitability also when tackling differing AR and ARMA processes and common factors, where the common AR($p$) restriction does not hold.
@@
In Supplement D.3, we compare the minimum eigenvalues of the predictors correlation matrix of ARMAr-LASSO with those of LASSO and GLS-LASSO. Results show that ARMAr-LASSO relies on a correlation matrix that exhibits a larger minimum eigenvalue than the classical LASSO and GLS-LASSO. Notably, this corroborates the statement of Remark 6.